40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sat Jun 8 00:24:11 2013, Annalisa, Update, Green Locking, Y arm locked with green but bad mode matching 
    Reply  Mon Jun 10 21:39:26 2013, Annalisa, Update, Green Locking, Y arm locked with green but bad mode matching 
       Reply  Tue Jun 11 10:00:54 2013, nicolas, Update, Green Locking, Y arm locked with green but bad mode matching 
          Reply  Wed Jun 12 01:56:58 2013, Annalisa, Update, Green Locking, Y arm locked with green but bad mode matching waistVaryingLensPos.pdfgreen.JPGModeMatchingY.pdf
             Reply  Wed Jun 12 22:48:10 2013, Annalisa, Update, Green Locking, Y arm locked with green but bad mode matching 
                Reply  Wed Jun 12 23:06:33 2013, Jenne, Update, Green Locking, Y arm locked with green but bad mode matching 
                   Reply  Thu Jun 13 00:35:23 2013, Manasa, Update, Green Locking, Y arm locked with green but bad mode matching 
                      Reply  Thu Jun 13 10:57:36 2013, Annalisa, Update, Green Locking, Y arm locked with green but bad mode matching 
                Reply  Thu Jun 13 16:01:48 2013, nicolas, Update, Green Locking, Y arm locked with green but bad mode matching 
                   Reply  Thu Jun 13 23:28:40 2013, Annalisa, Update, Green Locking, Y arm locked with green but bad mode matching 
Message ID: 8691     Entry time: Sat Jun 8 00:24:11 2013     Reply to this: 8692
Author: Annalisa 
Type: Update 
Category: Green Locking 
Subject: Y arm locked with green but bad mode matching 

[Annalisa, Nic]

After connecting the PD with the reflection from the arm  to the PDH box, theY  arm has been locked on the 01 mode.  Maximizing the alignment, we obtained a 00 mode locking, but we couldn't maximize the power.

The size of the reflected beam was different with respect to the size of the incoming beam, so probably a bad mode matching was one of the issues.

Moreover, the reflected beam is very low power. We need to figure out why it is so (bad alignment? related to mode matching?)

 

After measuring better all the distances, I did a new mode matching calculation. I put the lenses after measuring the beam waist, so the size of the beam on the lenses was the same as expected from the calculation. Nevertheless, the beam size on the beam splitter looks bigger than expected, and also in this case green flashes into the cavity at some HOM (again 01).

I also tried to lock again the cavity and maximize the alignment, but I didn't get any improvement with respect to the previous mode matching.

 

ELOG V3.1.3-